Are gender and age-associated with each other for emergency room visits? Do a chi-squared test for the following.   Please report the expected counts as a matrix. Also, is this a test of independence or not? Make the 4 steps cleared as outlined on page 325 of your textbook.

The following data represent the gender of 489 randomly selected patients who visited the emergency room with an injury-related emergency.

Gender (Age) Under 15 15-24 25-44 45-64 65-74 75 and older
Male 66 56 92 37 8 11
Female 44 39 66 34 12 24
Answer #1

Solution:

This is a test of independence.

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: Gender and age are not associated with each other for emergency room visits.

Alternative hypothesis: Ha: Gender and age are associated with each other for emergency room visits.

We assume level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 2

Number of columns = c = 6

Degrees of freedom = df = (r – 1)*(c – 1) = 1*5 = 5

α = 0.05

Critical value = 11.0705

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Age Group

Gender

<15

15 to 24

25 to 44

45 to 64

65 to 74

> 75

Total

Male

66

56

92

37

8

11

270

Female

44

39

66

34

12

24

219

Total

110

95

158

71

20

35

489

Expected Frequencies

Age Group

Gender

<15

15 to 24

25 to 44

45 to 64

65 to 74

> 75

Total

Male

60.7362

52.45399

87.23926

39.20245

11.04294

19.32515

270

Female

49.2638

42.54601

70.76074

31.79755

8.957055

15.67485

219

Total

110

95

158

71

20

35

489

Calculations

(O - E)

5.263804

3.546012

4.760736

-2.20245

-3.04294

-8.32515

-5.2638

-3.54601

-4.76074

2.202454

3.042945

8.325153

(O - E)^2/E

0.456196

0.239719

0.259798

0.123737

0.8385

3.586423

0.562434

0.295544

0.320299

0.152553

1.033768

4.421618

Chi square = ∑[(O – E)^2/E] = 12.29059

P-value = 0.031016

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

There is sufficient evidence to conclude that Gender and age are associated with each other for emergency room visits.